• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 00:25:50 on August19,2003

    Entry number 111867

    keyword=lost He gas flow for 10 min in septum

    the first fig shows the septum quench that was part of the training. at the far right of the strip (which was about 15 min ago), you can see a drop in the He flow (purple,pink lines) the LHe level also dropped after filling back up (orange line). the second fig show a zoom in on the last 1/2 hr. on this one, you can see the flow is back on? what happened? (this affected our spectrometers as well).

    i called MCC. they told me to call CHL. i called CHL. they said that they had another CHL trip at 1130, but that shouldn't affect us. they also said the the ESR looked okay. they don't sound like they know what happened.

    FIGURE 1

    FIGURE 2